# Optimal Transport-Based Decentralized Multi-Agent Distribution Matching

### v1

@@ I Introduction
Literature Survey: Coverage control represents one of the most established distributed approaches for organizing agents in space. It provides a classical framework in which agents optimize a sensing-based coverage functional. Foundational methods are based on Voronoi-partition gradients [cortes2004coverage], and subsequent work developed adaptive distributed implementations for networked robots [schwager2009decentralized]. Recent extensions address heterogeneous environments [liu2021multi], safety [bai2023safe], and online learning of unknown density fields [zhang2024multi]. While well suited for persistent sensing and exploration, coverage control converges to centroidal Voronoi configurations and therefore cannot enforce a prescribed terminal distribution.
@@
A complementary direction focuses on distribution steering using mean–field and continuum models. Zheng et al. [zheng2021transporting] proposed mean–field feedback laws, but these approaches rely on global density information and continuum approximations. Cui and Li [cui2024density] developed predictive density regulation with robustness guarantees, yet still assume centralized density availability and typically do not provide decentralized finite-agent, discrete-time controllers. Survey work in mean–field swarm theory [elamvazhuthi2019mean] highlights these structural assumptions, which limit applicability under local communication and agent-level constraints.
@@
Optimal Transport (OT)–based methods offer another viewpoint by modeling distribution evolution geometrically. Krishnan and Martínez [krishnan2022multiscale] interpreted coverage algorithms as gradient flows in Wasserstein space, and their recent work proposed a distributed online OT optimizer [krishnan2025distributed]. Although computationally scalable, these approaches operate at the optimization layer and do not directly yield decentralized feedback controllers for agents with discrete-time dynamics. Moreover, OT-based density regulation typically acts on continuous probability fields rather than explicit finite-agent terminal configurations [chen2016optimal, bandyopadhyay2014probabilistic].
@@
Non–uniform area coverage methods instead align time-averaged trajectories with a reference density map. Ergodic metrics provide a trajectory-level alignment tool [mathew2011metrics, rao2024learning], and Ivić et al. [ivic2019autonomous] demonstrated spatially adaptive spraying strategies. Recent density–driven optimal control (D2OC) frameworks steer time-averaged densities toward reference maps while respecting agent dynamics and physical constraints [seo2025smcs, seo2025tcst]. Such methods shape evolving densities but do not specify where agents ultimately reside at the mission’s end.
These research directions are fundamentally different from the terminal distribution matching problem considered here. Coverage and ergodic methods optimize long-term averages, and mean-field or OT-based density control operates at the continuum level rather than for finite agents. In contrast, our objective is to guide a finite team toward a prescribed discrete target distribution. This requirement arises in tasks such as positioning inspection robots at fixed infrastructure locations or arranging agents into a sampling grid for single-shot environmental measurements, where precise final placement rather than time-averaged behavior is essential.
@@ II Problem Formulation
Let the measures $\mu$ and $\nu$ be are supported on point sets $\{x_{i}\}_{i=1}^{M}$ and $\{y_{j}\}_{j=1}^{N}$, where $M$ denotes the number of agents contributing to the empirical distribution and $N$ denotes the number of target sample points representing the prescribed distribution $\nu$, with corresponding weights $\{\alpha_{i}\}$ and $\{\beta_{j}\}$, respectively. In this discrete setting, the optimal transport problem between $\mu$ and $\nu$ reduces to computing the squared 2-Wasserstein distance (also known as the Kantorovich formulation [villani2008optimal]),
@@
In the discrete formulation, the coefficients $\alpha_{i}$ represent the mass or importance assigned to each agent location $x_{i}$. Throughout this work we assume $\alpha_{i}=1/M$, reflecting that each agent contributes equally to the empirical distribution. However, non-uniform values are also admissible. For example, if certain agents possess higher sensing resolution, stronger actuation capability, or play a more influential role in the team’s decision-making hierarchy, one may assign larger weights to the corresponding $\alpha_{i}$. In contrast, the coefficients $\beta_{j}$ specify the mass prescribed by the target distribution at each sample location $y_{j}$ and encode the structure of the desired reference density.
@@
where $\mu_{k}=\frac{1}{M}\sum_{i=1}^{M}\delta_{x_{i}(k)}$ is empirical distribution of the agent states at time $k$, and $\nu=\frac{1}{N}\sum_{j=1}^{N}\delta_{y_{j}}$ represents the reference distribution supported on the target points $\{y_{j}\}_{j=1}^{N}$. Because each agent state $x_{i}(k)$ evolves over time under the system dynamics, the empirical distribution $\mu_{k}$ is time-varying. As a consequence, the objective $\mathcal{W}_{2}(\mu_{k},\nu)$ is also a time-dependent quantity evaluated along the closed-loop evolution of the multi-agent system.
@@ III-A Sample Selection Phase: Sequential Target Assignment
    Solve Local Transport Problem: Given a single agent at $x_{i}(k)$, the local OT problem admits an analytic greedy solution. Define the distances $d_{j}:=\|x_{i}(k)-y_{j}\|$ and reorder the indices so that $d_{j_{1}}\leq d_{j_{2}}\leq\cdots$. The agent then allocates portions of its mass $1/M$ to the nearest sample points in this order. Specifically, $\pi_{ij_{\ell}}=\min\!\left(\beta_{j_{\ell}},\;\frac{1}{M}-\sum_{m<\ell}\pi_{ij_{m}}\right)$, where the term $\frac{1}{M}-\sum_{m<\ell}\pi_{ij_{m}}$ represents the remaining mass after allocations to the closer sample points $j_{1},\dots,j_{\ell-1}$. This greedy nearest–neighbor allocation continues until the full mass $1/M$ is exhausted and yields a closed–form local transport plan.
@@ IV-A Linear Time-Invariant System Case
The result follows immediately from the Moore–Penrose pseudoinverse characterization of minimum-norm solutions. Hence, the proof is omitted.
@@ IV-B Control-Affine Nonlinear System Case
For the control–affine dynamics in (3) and the finite–horizon cost (13), the necessary conditions for optimality follow directly from Pontryagin’s minimum principle. The resulting state–costate recursion and optimal control law are:
@@
where $\omega_{i}=\sum_{j\in\mathcal{S}_{i}}\tilde{\pi}_{ij}$ and $y_{i}^{*}$ is the weighted barycenter of the assigned sample points. These relations form the two–point boundary value system used to compute the nonlinear control input.
The derivation of these optimality conditions is not included, since they follow directly from the standard Pontryagin minimum principle for discrete–time control–affine systems. We summarize the result here only to state the conditions required for computing the nonlinear controller.
@@ V-C Weight Update with Memory Decay
The decay factor $\gamma$ adjusts how strongly outdated information is suppressed: $\gamma=0$ disables memory correction, while larger values subtract more previously stored information from the current weights, reducing sample-point overlap under intermittent communication. The element-wise minimum ensures a conservative adjustment, compensating for outdated information while avoiding overrepresentation of already-covered areas. The element-wise minimum ensures a conservative adjustment, compensating for outdated information while avoiding overrepresentation of already-covered areas. Finally, an element-wise non-negativity projection is applied to ensure $\boldsymbol{\underline{\beta}}_{i}^{(k)}\succeq 0$ as follows: $\boldsymbol{\underline{\beta}}_{i}^{(k)}\leftarrow\max\big(\boldsymbol{0},\boldsymbol{\underline{\beta}}_{i}^{(k)}\big)$ (element-wise).
@@ VI Cycle-Level Convergence Guarantees
At the beginning of each cycle, every agent commits to a fixed local transport plan $\tilde{\pi}_{ij}$ determined by the preceding selection phase and follows the corresponding control inputs during the cycle. This section shows that the resulting evolution possesses a monotone cycle-wise descent property and establishes a cycle-level bound on the discrepancy between the agent distribution and the target distribution. These results provide the key convergence guarantees of the proposed framework.
@@
Evaluating (23) at the beginning and end of the cycle, we obtain
@@ VII-A LTI System Case
Four scenarios are considered: Figs. 1(a), (c), and (d) use a concentrated initial distribution $x(0)$, uniformly sampled from a small square and Fig. 1(b) uses a uniform distribution across the entire domain.

### v2

@@ I Introduction
Literature Survey: Coverage control represents one of the most established distributed approaches for organizing agents in space. It provides a classical framework in which agents optimize a sensing-based coverage functional. Foundational methods are based on Voronoi-partition gradients [4], and subsequent work developed adaptive distributed implementations for networked robots [13]. Recent extensions address heterogeneous environments [10], safety [1], and online learning of unknown density fields [17]. While well-suited for persistent sensing and exploration, coverage control converges to centroidal Voronoi configurations and therefore cannot enforce a prescribed terminal distribution.
@@
A complementary direction focuses on distribution steering using mean-field and continuum models. Zheng et al. [18] proposed mean-field feedback laws, but these approaches rely on global density information and continuum approximations. Cui and Li [5] developed predictive density regulation with robustness guarantees, yet still assume centralized density availability and typically do not provide decentralized finite-agent, discrete-time controllers. Survey work in mean-field swarm theory [6] highlights these structural assumptions, which limit applicability under local communication and agent-level constraints.
@@
Optimal Transport (OT)-based methods offer another viewpoint by modeling distribution evolution geometrically. The authors in [8] interpreted coverage algorithms as gradient flows in Wasserstein space, and their recent work proposed a distributed online OT optimizer [9]. Although computationally scalable, these approaches operate at the optimization layer and do not directly yield decentralized feedback controllers for agents with discrete-time dynamics. Moreover, OT-based density regulation typically acts on continuous probability fields rather than explicit finite-agent terminal configurations [3, 2].
@@
Nonuniform area coverage methods instead align time-averaged trajectories with a reference density map. Ergodic metrics provide a trajectory-level alignment tool [11, 12], and Ivić et al. [7] demonstrated spatially adaptive spraying strategies. Recent density-driven optimal control (D2OC) frameworks steer time-averaged densities toward reference maps while respecting agent dynamics and physical constraints [15, 14]. Such methods shape evolving densities but do not specify where agents ultimately reside at the mission’s end.
While these research directions share the goal of shaping collective behavior, our approach focuses on terminal distribution matching for discrete agents. Unlike ergodicity-based methods or $D^{2}OC$ that optimize time-averaged trajectories, we explicitly determine finite-agent trajectories to satisfy a prescribed final configuration. This work is the first to address terminal matching within discrete agent systems, providing a clear departure from continuum-level PDE steering or trajectory alignment over time. This distinction is vital for tasks like establishing static agent formations for monitoring or coordinated multi-robot inspection, where the final static placement of each agent is more critical than its cumulative coverage behavior.
@@ II Problem Formulation
Let the measures $\mu$ and $\nu$ be supported on point sets $\{x_{i}\}_{i=1}^{M}$ and $\{y_{j}\}_{j=1}^{N}$, where $M$ denotes the number of agents contributing to the empirical distribution and $N$ denotes the number of target sample points representing the prescribed distribution $\nu$, with corresponding weights $\{\alpha_{i}\}$ and $\{\beta_{j}\}$, respectively. In this discrete setting, the optimal transport problem between $\mu$ and $\nu$ reduces to computing the squared 2-Wasserstein distance (also known as the Kantorovich formulation [16]),
@@
In the discrete formulation, the coefficients $\alpha_{i}$ represent the mass or importance assigned to each agent location $x_{i}$. Throughout this work, we assume $\alpha_{i}=1/M$, reflecting that each agent contributes equally to the empirical distribution. However, non-uniform values are also admissible. For example, if certain agents possess higher sensing resolution, stronger actuation capability, or play a more influential role in the team’s decision-making hierarchy, one may assign larger weights to the corresponding $\alpha_{i}$. In contrast, the coefficients $\beta_{j}$ specify the mass prescribed by the target distribution at each sample location $y_{j}$ and encode the structure of the desired reference density.
@@
where $\mu_{k}=\frac{1}{M}\sum_{i=1}^{M}\delta_{x_{i}(k)}$ is the empirical distribution of the agent states at time $k$, and $\nu=\frac{1}{N}\sum_{j=1}^{N}\delta_{y_{j}}$ represents the reference distribution supported on the target points $\{y_{j}\}_{j=1}^{N}$. Because each agent state $x_{i}(k)$ evolves over time under the system dynamics, the empirical distribution $\mu_{k}$ is time-varying. As a consequence, the objective $\mathcal{W}_{2}(\mu_{k},\nu)$ is also a time-dependent quantity evaluated along the closed-loop evolution of the multi-agent system.
@@ III-A Sample Selection Phase: Sequential Target Assignment
    Solve Local Transport Problem: Given a single agent at $x_{i}(k)$, the local OT problem admits an analytic, greedy solution. Define the distances $d_{j}:=\|x_{i}(k)-y_{j}\|$ and reorder the indices so that $d_{j_{1}}\leq d_{j_{2}}\leq\cdots$. The agent then allocates portions of its mass $1/M$ to the nearest sample points in this order. Specifically, $\pi_{ij_{\ell}}=\min\!\left(\beta_{j_{\ell}},\;\frac{1}{M}-\sum_{m<\ell}\pi_{ij_{m}}\right)$, where the term $\frac{1}{M}-\sum_{m<\ell}\pi_{ij_{m}}$ represents the remaining mass after allocations to the closer sample points $j_{1},\dots,j_{\ell-1}$. This greedy nearest-neighbor allocation continues until the full mass $1/M$ is exhausted and yields a closed-form local transport plan.
@@ IV-A Linear Time-Invariant System Case
The result follows immediately from the Moore-Penrose pseudoinverse characterization of minimum-norm solutions. Hence, the proof is omitted.
@@ IV-B Control-Affine Nonlinear System Case
For the control-affine dynamics in (3) and the finite-horizon cost (13), the necessary conditions for optimality follow directly from Pontryagin’s minimum principle. The resulting state-costate recursion and optimal control law are:
@@
where $\omega_{i}=\sum_{j\in\mathcal{S}_{i}}\tilde{\pi}_{ij}$ and $y_{i}^{*}$ is the weighted barycenter of the assigned sample points. These relations form the two-point boundary value system used to compute the nonlinear control input.
The derivation of these optimality conditions is not included, since they follow directly from the standard Pontryagin minimum principle for discrete-time, control-affine systems. We summarize the result here only to state the conditions required for computing the nonlinear controller.
@@ V-C Weight Update with Memory Decay
The decay factor $\gamma$ adjusts how strongly outdated information is suppressed: $\gamma=0$ disables memory correction, while larger values subtract more previously stored information from the current weights, reducing sample-point overlap under intermittent communication. The element-wise minimum ensures a conservative adjustment, compensating for outdated information while avoiding overrepresentation of already-covered areas. Finally, an element-wise non-negativity projection is applied to ensure $\boldsymbol{\underline{\beta}}_{i}^{(k)}\succeq 0$ as follows: $\boldsymbol{\underline{\beta}}_{i}^{(k)}\leftarrow\max\big(\boldsymbol{0},\boldsymbol{\underline{\beta}}_{i}^{(k)}\big)$ (element-wise).
@@ VI Cycle-Level Convergence Guarantees
At the beginning of each cycle, every agent commits to a fixed local transport plan $\tilde{\pi}_{ij}$ determined by the preceding selection phase and follows the corresponding control inputs during the cycle. This section shows that the resulting evolution possesses a monotonic, cycle-wise descent property and establishes a cycle-level bound on the discrepancy between the agent distribution and the target distribution. These results provide the key convergence guarantees of the proposed framework.
@@
Evaluating (23) at the beginning and the end of the cycle, we obtain
@@ VII-A LTI System Case
Two scenarios are considered: Fig. 1(a) shows a concentrated initial distribution $x(0)$, uniformly sampled from a small square while Fig. 1(b) displays a uniform distribution across the entire domain.